|  September 2, 2014  |  
Partly Cloudy   66.0F  |  Forecast »
Bookmark and Share Email this page Email Print this page Print Feed Feed

Repsol Nabors Qugruk 2 UPDATE Crews continue steam thawing and removing drilling mudQ2 Pad Gas and Mud Release SITREP # 10

ALASKA DEPARTMENT OF ENVIRONMENTAL CONSERVATION
Division of Spill Prevention and Response
Prevention and Emergency Response Program

SITUATION REPORT

Repsol Q2 Pad Gas and Mud Release

SITREP # 10

SPILL NUMBER: 12399904601

TIME/DATE OF SPILL: Repsol reported the incident to ADEC at 10:29 AM on February 15, 2012.

TIME/DATE OF SITUATION REPORT: 4:00 PM on February 23, 2012

TYPE/AMOUNT OF PRODUCT SPILLED: Repsol estimates that approximately 42,000 gallons (1,000 barrels) of
fresh‐water‐based drilling mud have been released to the ice drilling pad and adjacent snow‐covered tundra. An
unknown amount of gas has been released. The mud that was released is a mixture of minerals (primarily
barium sulfate) and water. Drilling mud is used to lubricate the drill, control formation pressure, and remove
cuttings from the bore. No oil has been spilled.

INCIDENT LOCATION: Qugruk #2 pad (Q2 pad), on the Colville River Delta, approximately 18 miles northeast of
Nuiqsut and approximately 150 miles southeast of Barrow (70° 27’ 19” N, 150° 44’ 52” W).

CAUSE OF SPILL: Repsol contractor Nabors Drilling was drilling an exploratory well when the drill penetrated a
shallow gas pocket at a depth of 2,523 feet, resulting in a gas kick. The gas kick drove drilling mud out of the
well and through the gas diverter onto the ice pad and adjacent snow‐covered tundra. The gas diverter is a pipe
that extends approximately 75 feet from the drilling rig toward the edge of the pad. Additional mud was
pumped into the well in an attempt to control it, but that mud was also forced out by the gas.

POTENTIAL RESPONSIBLE PARTY (PRP): Repsol E&P USA

RESPONSE ACTION: Response crews continue twenty‐four hour operations. Two large heaters pump hot air into
the drilling rig for thawing necessary equipment. Crews continue steam thawing and removing drilling mud.
Removal remains focused on the waste mud material that blocks access to essential drilling rig components
needed for well‐kill operations. Complete drilling rig cleanup will occur after the well has been successfully
killed and secured.

An inspection camera was received onsite and used to evaluate the bore hole integrity, which appears to be in
good condition. This will allow the pumping of drilling mud down the well for the final step in the well‐kill
process. Meanwhile, the hydraulic system was function‐tested and no leaks were observed.

Repsol reports that 671 bbls of liquids (thawed drilling mud and water) have been manifested and shipped
offsite for disposal. Approximately 340 cubic yards of frozen drilling mud waste material is stockpiled on a
portion of the ice pad and awaiting transport by trucks for disposal at the designated grind and inject facility.

SOURCE CONTROL: The well ceased flowing at about 9:00 PM on February 16. Preparations for a well‐control
operation are ongoing (see above).

RESOURCES AFFECTED: Snow‐covered tundra and ice pad. Drilling mud can affect tundra plants by changing
soil pH and salinity. Brackish water produced by the well may also increase soil salinity.

FUTURE PLANS AND RECOMMENDATIONS: Cleanup cannot safely begin until the well is under control. The
well‐control team will give cleanup personnel approximately 24‐hours notice in advance of completion of the
well‐control operation. Plans are in place for delineating and cleaning up the spill and necessary personnel
crews and equipment are on standby. A reduced incident management team will continue to monitor the
operation from Anchorage and two ADEC responders will remain at Palm Pad until cleanup operations are
imminent, at which time additional staff will be recalled.

UNIFIED COMMAND AND PERSONNEL:
Incident Commander: Jeremy Michels, O’Brien’s Response Management
F.O.S.C.: Matt Carr, EPA
S.O.S.C.: Tom DeRuyter, ADEC
L.O.S.C.: Gordon Brower, North Slope Borough
Field SOSC: Wes Ghormley, ADEC

WEATHER: Tonight: Periods of snow tapering off to flurries, areas of fog; less than one inch snow
accumulation; lows around 22 below; variable winds less than 15 mph.
Friday: Cloudy, flurries and patchy fog; highs around 15 below; variable winds less than 15 mph.

TIME/DATE OF THE NEXT SITUATION REPORT: 4:00 PM on February 24, 2012

FOR ADDITIONAL INFORMATION CONTACT: Ty Keltner, Public Information Officer, ADEC (907) 465‐5009